Two pipes A and B, are used to fill a water tank. The empty tank is filled in 10 hours if two pipes are used together. If pipe A alone is used for 6 hours and then turned off, pipe B will .take over and finish off filling the tank in 18 hours. How long will it take each pipe alone to fill the tank?
Answer by htmentor(1343)

Let A = the fill rate of pipe A
Let B = the fill rate of pipe B
Working together, their combined rate is A + B = 1 tank/10 h = 1/10 -> A = 1/10 - B
Pipe A working alone will fill A*6 of a tank
The remaining portion of the tank is filled by B in 18 h
Thus 1 = A*6 + B*18
1 = 6(1/10 - B) + 18B
Solving for B gives 12B = 4/10 -> B = 1/30
Thus A = 1/10 - 1/30 = 2/30
So it would take A 15 hrs and B 30 hrs to fill the tank alone